OpenStudy (faiqraees):

A very confusing question How would the mathematical expression 6 / 2(1+2) would be expressed \[\large\rm \frac{6}{2} * (1+2)\] Argument: This is the correct expression because the given question is 6/2 * (2+1) and thus division should be done first giving the answer 9 OR \[\large\rm \frac{6}{2(1+2)}\] Argument: Multiplication should be done first because 2 is being multiplied with a bracket giving the answer 1

OpenStudy (faiqraees):

The expression is \[\large\rm 6\div2(1+2)\]

hartnn (hartnn):

6/2(1+2) is always 6/2*3 = 3*3 << Important step = 9

OpenStudy (faiqraees):

However calculators give the answer 1

hartnn (hartnn):

why? when we have equal priority operation, we evaluate from \(\Large \text{left to right}.\)

hartnn (hartnn):

online calculators are designed to work the proper way. http://www.wolframalpha.com/input/?i=6%2F2(1%2B2) the calculator you have might be designed to work from right to left.

OpenStudy (wcrmelissa2001):

my calculator gives the answer 9...is it an old model?

OpenStudy (faiqraees):

@hartnn what if the question was 6/2x would you write it 3x?

OpenStudy (faiqraees):

@ganeshie8 @mathmale

hartnn (hartnn):

absolutely 6/2x is 3x 6/(2x) is 3/x